OpenStudy (anonymous):

Two similar solid statues are 20 cm and 30 cm tall. If the shorter one weighs 6 kg, how much will the taller one weigh?

OpenStudy (anonymous):

weigh = volume volume of A : volume of B = (length of A) ^3 : (length of B)^3 6:x=20^3:30^3 6/x = 20^3 / 30^3 solve this to find x

OpenStudy (p0sitr0n):

\[6/x =(2/3)^3\] \[6/(2/3)^3\] =x x =20.25

OpenStudy (anonymous):

how did you know to put 8000/27000?

OpenStudy (anonymous):

instead of 27,000/8,000

OpenStudy (p0sitr0n):

yeah true, but you can easuly reverse it and get 9kg as answer

OpenStudy (anonymous):

oh so it could be either one?

OpenStudy (anonymous):

my book has 20.25

OpenStudy (anonymous):

take a cube as an example, (length of A)^3 = volume of A (length of B)^3 = volume of B volume of A/ volume of B = (length of A)^3 / (length of B)^3 so, it cannot be reversed!